OpenStudy (anonymous):

Chantel is carpeting a room that has an area of x^2-100 square feet. If the width of the room is x-10 feet, what is the length of the room.

OpenStudy (anonymous):

Here's what i've done so far: -x+10+x^2-100= l(x-10)-x+10 (-x+10)+(x^2-100)=l x^2-x-90=l WHAT NEXT?

OpenStudy (anonymous):

I don't understand your work at all. All you have to do is divide the area by the width.

OpenStudy (anonymous):

so that would just give me the length?

hartnn (hartnn):

yes, just divide (x^2-100) by x-10 can u divide it?

OpenStudy (anonymous):

length x width = area is it not? 5th grade stuff.

OpenStudy (anonymous):

ok. one sec

OpenStudy (anonymous):

but 100/10 is 10. but then how do i divide x^2 by x?

OpenStudy (anonymous):

exponent laws, what happens when you divide identical bases with different exponents?

hartnn (hartnn):

thats not the way to divide polynomials.... wait i will give you a better approach than dividing like that.... do u know the formula: \[a ^{2}-b^{2}=(a+b)(a-b)\]

OpenStudy (anonymous):

yeah

hartnn (hartnn):

so u can write this : \[x^2-100 = x^2-10^2 = (x+10)(x-10)= AREA\] got this?

OpenStudy (anonymous):

right

hartnn (hartnn):

now using this : AREA=WIDTH *LENGTH,can i write : \[(x+10)(x-10)= (x-10) *LENGTH\] so what will be lenght??

OpenStudy (anonymous):

ok. one sec

sam (.sam.):

You know that Area=Length x Height (x^2-100)=(x-10)( Height ) \[(Height)=\frac{(x^2-100)}{(x-10)}\] Factoring \(x^2-100\), \[(Height)=\frac{(x+10)(x-10)}{(x-10)}\] \[(Height)=\frac{(x+10)\cancel{(x-10)}}{\cancel{(x-10)}}\] \[(Height)=x+10\]

OpenStudy (anonymous):

hartnn, to solve (x+10)(x-10) i use FOIL, right. i got x^2-100.

sam (.sam.):

@Amenah , yes because Area= (x+10)(x-10) Area=x^2-100 So you're right

hartnn (hartnn):

see @.Sam. comment,it explains everything in one place

OpenStudy (anonymous):

k

hartnn (hartnn):

ask if u still have doubts.....

OpenStudy (anonymous):

for the length, i got x+10

hartnn (hartnn):

yes,correct...thats your answer....understood all steps clearly?

OpenStudy (anonymous):

yep. thank you so much! i had so many question on this packet, and now i'm almost done! can i ask you one more question?

hartnn (hartnn):

ask as many as u want....